OpenStudy (anonymous):

solve by substitution; 2x+2y=14 y=3x-9 please/thanks:)

OpenStudy (anonymous):

easy again bro. just substitute y=3x-9 into 2x+2y=14 to get 2x + 2(3x-9)=14 then you get 2x + 6x -18=14. then you get 8x=32, which means x=4. then just put x=4 into y=3x-9 to get y=3.

OpenStudy (anonymous):

THANKS! i get it now, i was absent from school and missed the lesson/:

OpenStudy (anonymous):

no prob man!

OpenStudy (anonymous):

Wait, but what about this one; -7x=2y=0 x-6y=0